Mr.Ali income elasticity of meat demand is 3. How much his demand for meat will increase when his income increases from Rs 500 to Rs 1000. Is meat normal or Inferior good for him?.
1
Expert's answer
2020-08-20T10:18:08-0400

Solution:

"Income\;elasticity \;of\;demand= \frac{\%\;change\;in\;demand }{\%\;change\;in\;income}"


"3=\frac{\%\;change\;in\;demand }{(\frac{1000-500}{500} ) }"


"3=\frac{\%\;change\;in\;demand}{100\%}"


"3\times 100 \% = \% \; change\; in\; demand"


"\%\;change\;in\;demand=300\%"


Therefore the new meat demand is:

"300\%\times 3=9"

Hence his demand for meat increased by:

"9-3=6"


The meat is a normal good for him

This is because a normal good is any good in which demand increases when income increases, and the demand for meat has substantially increased with the increase in income levels.